# Neural Object Detection for 4D STEM: High-Throughput Sub-Pixel Electron Diffraction Pattern Recognition

### v1

@@ 1 Introduction
Figure 1: Top left to right: diagram of the neural diffraction pattern detection (NDPD) network along with exemplary activation maps from the final C2f layer, showcasing kernel responses to an electron diffraction pattern at stride levels of 8, 16, and 32. Bottom left to right: rapid phase identification of a complex phase transformed Ti-50Nb metallic alloy by the network.
Transmission electron microscopy (TEM) has emerged as an essential technique in materials science, functioning as a versatile tool for both academic and industrial applications [41]. In scanning TEM (STEM), a focused electron probe, typically with a sub-angstrom full-width half maximum (FWHM), is scanned across a thin region of interest (ROI). The interaction of the electron beam with the material results in both elastic and inelastic scattering signals. These signals are gathered by various angular imaging detectors, offering immediate feedback on the sample’s morphology and crystallography.
@@
In 2D-STEM, a stationary diffraction pattern inherently forms during imaging; however, a diffraction pattern for every probe position is not recorded. The introduction of 4D-STEM bridges this gap by enabling the acquisition and storage of each diffraction pattern using high-speed detectors, such as complementary metal-oxide-semiconductor (CMOS) or electron-pixel array detectors [44, 4].
@@
Historically, electron diffraction patterns have been analyzed either through measurement from digitally recorded images or via semi-automatically using sequential pipelines involving peak-localization algorithms, such as 2D Gaussian fitting, peak pair analysis, minimum enclosing circle, center of mass calculations, Hough transform [7, 42, 2, 51, 5] These algorithms calculate Bragg diffraction spacings in reciprocal space or in real space using Fourier transforms. It is well-established that the precision of the position calculations by these pipelines is highly sensitive to the signal-to-noise ratio (SNR) of electron diffraction patterns, dynamical effects, variations in sample tilt and thickness [20, 34, 32, 2, 5].
@@
Fig. 1 displays the main components of the neural diffraction pattern detection (NDPD) architecture [35, 18] and kernel activation responses from the network’s final C2f layers. Accompanying this is a benchmark example, a phase map generated by the network from a lamellar, multiphase microstructure Ti-50Nb metallic alloy. The phase identification map reveals lamellar regions of the $\alpha$-hcp (hexagonal close-packed) phase and the $\beta$-bcc (body-centered cubic) matrix, along with overlapping zones where the lamellar phase intersects with the matrix.
@@ 2 Related work
Recently, cross-correlation methods that reference a standard diffraction disc or probe image template have been adopted for the analysis of convergent beam electron diffraction (CBED) patterns with large diffraction discs [34, 9, 8, 30, 47, 25]. These methods mitigate post-processing ambiguities stemming from dynamical effects in CBED patterns. Cross-correlation has proven particularly effective when using direct electron detectors with limited dynamic range, where electron dose must be regulated by distributing the CBED discs across a large number of pixels in the detector array [34].
@@
Machine learning (ML) has emerged as a powerful tool for the processing of 4D-STEM datasets [21, 52, 39, 27]. Yoo et al [50] introduced a methodology that combines dimensionality reduction, cepstral analysis, and unsupervised ML via k-means clustering to extract interatomic distances from complex metallic microstructures. Transformation of diffraction patterns into the real-space cepstral domain effectively isolates coexisting phases and maps strain fields in a metallic shape-memory alloy.
In the proposed pipeline taxonomy, unsupervised algorithms like k-means clustering entail nonlinear computations, such as minimizing Euclidean distance. However, they are often executed in deterministic, sequential methods [3, 50], which makes them structurally similar to conventional image processing approaches, unlike the dynamic, end-to-end framework of modern deep learning models.
@@
However, CNN architectures, such as U-Net, often exhibit limited generalization, particularly with respect to variations in feature scale [37], with model performance closely tied to the quantity and diversity of the training data [10, 11, 13, 15, 46, 43]. Moreover, while simulated datasets are valuable for controlled training, they are often simplistic and may fail to capture the full complexity of the experimental conditions [43, 46, 11].
@@ 3.1 Bragg disc detection and localization
Single-stage object detection models, such as YOLOv8, adopt a fully convolutional framework that directly predicts bounding box coordinates from feature maps [18, 45]. This anchor-free design enables a well-generalized [11], object-center-aware feature augmentation, where grid cells serve as reference points for predicting offsets to object centers. Such a center-based approach is particularly well-suited for the high-throughput and high-precision detection of Bragg disc centers in electron diffraction patterns [35, 18, 6, 45].
@@
Where $c_{x}$ and $c_{y}$ are the column and row indices of the grid cell, and $s$ is the stride level.
To improve the center localization in our object detection model, we incorporated an efficient autocorrelation-based approach, leveraging the Wiener–Khinchin theorem [48, 19]. Recently, the log-magnitude power spectrum of reciprocal-space intensities has been applied via cepstral analysis to decouple lattice periodicities in 4D-STEM diffraction patterns. Although the logarithmic operation compresses the dynamic range and suppresses dominant intensity variations, it can potentially obscure weaker signals in low SNR situations. In our implementation, spectral decomposition is performed without applying the logarithmic transform, similar to the Patterson function [33], utilizing an autocorrelation function to retain the full-intensity scaling. This strategy enhances the extraction of interatomic distances while minimizing potential ambiguities in Bragg disc intensities caused by sample tilt and thickness variations.
@@ 4.2 Strain mapping of pseudomorphic Si/SiGe epitaxial layers
To investigate the fidelity of our NDPD workflow in strain analysis, we acquired a 4D-STEM dataset from a Si/SiGe multilayer TEM sample at an operating voltage of 200 kV. The acquisition conditions were optimized so that each diffraction disc spans at least 10 pixels in diameter and extends to at least the third-order Laue zone, ensuring both spatial precision and sufficient sampling of the diffraction space.
Fig. 6 illustrates a set of diffraction patterns from the Si substrate of the Si/SiGe TEM sample. The object detection model generates bounding box predictions, and from these predictions, we extract the center coordinates for each diffraction disc. The alternating Si and SiGe layers provide distinct light and dark contrast, as shown in the virtual annular dark-field STEM image in Fig. 7. Strain maps $\varepsilon_{xx}$ and $\varepsilon_{yy}$ represent the out-of-plane and in-plane strain fields populated from two orthogonal reciprocal lattice vectors, [100] and [110], respectively.
The strain calculations are conducted by tracking the center coordinates of diffraction discs by our object detection model and relative displacements measured with respect to a reference region selected in the strain-free Si substrate. The oscillations in the colored map of strain field $\varepsilon_{xx}$ indicate positive (tensile) strain along the growth direction. The profile is plotted along the [100] direction, and the corresponding mean strain variation is approximately 0.8 %, which is in good agreement with the expected maximum theoretical value in SiGe layers.
@@
The strain variations due to the bending of the thin TEM foil are also evident in the strain field maps. These variations contribute to the strain modulations on the order of 0.1-0.2 % along the nominally strain-free [110] direction in the $\varepsilon_{yy}$ strain map.
Computational benchmarking of the end-to-end latency for the strain analysis was performed using a 4D-STEM dataset acquired from an ROI scanned over a 97$\times$40 pixels field of view. A total of 3880 diffraction patterns, each with an image resolution of 512$\times$512 pixels, were processed asynchronously in 35 seconds, yielding an effective throughput of 110 fps. The model inference was carried out at an input resolution of 512$\times$512 pixels per pattern.
Figure 7: Top left to right: the virtual DF image of Si/SiGe multilayers, the uniaxial strain map $\varepsilon_{xx}$ along the growth direction, and the orthogonal strain component $\varepsilon_{yy}$. Bottom left to right: the line profile across the thin-film growth [100] and strain-free [110] directions.
@@ 4.3 Phase and strain mapping of a multiphase metallic alloy
Figure 8: Top left to right: example electron diffraction patterns for matrix/coherent Phase I and incoherent Phases II and III overlapping with bcc matrix. Middle left to right: virtual DF image, and interatomic lattice displacement map. Bottom left to right: strain maps of $\varepsilon_{xx}$ and $\varepsilon_{yy}$.
Fig. 8 presents example displacement field maps derived from two orthogonal [110] reciprocal lattice vectors, respectively. These maps reveal small, spherical bcc-precipitates (Phase I), approximately 50 nm in size, that are coherently embedded within the parent $\beta$-bcc matrix. In contrast, incoherent phases are identified as Phase II, corresponding to the ordered oxide $Ti_{3}O$, and Phase III, comprising hcp-precipitates that overlap with the parent bcc-matrix, as evidenced by the electron diffraction patterns. The corresponding strain maps reveal a tensile $\varepsilon_{xx}$ and compressive $\varepsilon_{yy}$ coherency strain fields originating from the embedded bcc-precipitates. These findings support strain engineering strategies that can be leveraged to improve the mechanical performance of the alloy.
@@ 5 Conclusion
In this work, we introduce a parallelized post-processing framework for large-scale 4D-STEM datasets, leveraging an efficient neural network-based object detection model. Rapid and accurate object detection presents a promising avenue for the analysis of electron diffraction patterns and the localization of interatomic lattice distances. Our approach enables sub-pixel precision in the detection of diffraction features with high precision and accuracy. The ability to rapidly and reliably extract crystallographic information from 4D-STEM datasets acquired under varying experimental conditions accelerates materials discovery, from semiconductor development to alloy design.

### v2

# Neural Object Detection for 4D-STEM: High-Throughput Sub-Pixel Electron Diffraction Pattern Recognition

@@ 1 Introduction
Figure 1: Top left to right: diagram of the neural diffraction pattern detection (NDPD) network along with exemplary activation maps from the final C2f layer, showcasing kernel responses to an electron diffraction pattern at stride levels of 8, 16, and 32. Bottom left to right: rapid phase identification of a complex phase-transformed Ti-50Nb metallic alloy by the network.
Transmission electron microscopy (TEM) has emerged as an essential technique in materials science, functioning as a versatile tool for both academic and industrial applications [48]. In scanning transmission electron microscopy (STEM), a focused electron probe, typically with a sub-angstrom full-width half maximum (FWHM), is scanned across a thin region of interest (ROI). The interaction of the electron beam with the material results in both elastic and inelastic scattering signals. These signals are gathered by various angular imaging detectors, offering immediate feedback on the sample’s morphology and crystallography.
@@
In two-dimensional (2D)-STEM, a stationary diffraction pattern inherently forms during imaging; however, a diffraction pattern for every probe position is not recorded. The introduction of four-dimensional (4D)-STEM bridges this gap by enabling the acquisition and storage of each diffraction pattern using high-speed detectors, such as complementary metal-oxide-semiconductor (CMOS) or electron-pixel array detectors [52, 5].
@@
Historically, electron diffraction patterns have been analyzed either through measurement from digitally recorded images or via semi-automatically using sequential pipelines involving peak-localization algorithms, such as 2D Gaussian fitting, peak pairs analysis, minimum enclosing circle, center of mass calculations, Hough transform [9, 49, 3, 60, 6, 32, 30, 7] These algorithms calculate Bragg diffraction spacings in reciprocal space or in real space using a Fourier transform. It is well-established that the precision of the position calculations by these pipelines is highly sensitive to the signal-to-noise ratio (SNR) of electron diffraction patterns, dynamical effects, variations in sample tilt and thickness [24, 41, 39, 3, 6].
@@
Fig. 1 displays the main components of the neural diffraction pattern detection (NDPD) architecture [42, 22] and kernel activation responses from the network’s final C2f layers. Accompanying this is a benchmark example, a phase map generated by the network from a lamellar, multiphase microstructure Ti-50Nb metallic alloy. The phase identification map reveals lamellar regions of the $\alpha$-hcp (hexagonal close-packed) phase and the $\beta$-bcc (body-centered cubic) matrix, along with overlapping zones where the lamellar phase intersects with the matrix. NDPD detects these overlapping regions by identifying double diffraction discs from two superimposed crystal structures.
@@ 2 Related work
Recently, cross-correlation methods that reference a standard diffraction disc or probe image template have been adopted for the analysis of convergent beam electron diffraction (CBED) patterns with large diffraction discs [34, 41, 12, 11, 37, 56, 29, 62]. These methods mitigate post-processing ambiguities stemming from dynamical effects in CBED patterns. Cross-correlation has proven particularly effective when using direct electron detectors with limited dynamic range, where electron dose must be regulated by distributing the CBED discs across a large number of pixels in the detector array [41, 35].
@@
Machine learning (ML) has emerged as a powerful tool for the processing of 4D-STEM datasets [25, 61, 46, 33, 10, 64, 53]. Yoo et al [59] introduced a methodology that combines dimensionality reduction, cepstral analysis, and unsupervised ML via k-means clustering to extract interatomic distances from complex metallic microstructures. Transformation of diffraction patterns into the real-space cepstral domain effectively isolates coexisting phases and maps strain fields in a metallic shape-memory alloy.
In the proposed pipeline taxonomy, unsupervised algorithms like k-means clustering entail nonlinear computations, such as minimizing Euclidean distance. However, they are often executed in deterministic, sequential methods [4, 59, 1, 50, 53], which makes them structurally similar to conventional image processing approaches, unlike the dynamic, end-to-end framework of modern deep learning models.
@@
However, CNN architectures, such as U-Net, often exhibit limited generalization, particularly with respect to variations in feature scale [44], with model performance closely tied to the quantity and diversity of the training data [13, 14, 16, 19, 55, 51, 61]. Moreover, while simulated datasets are valuable for controlled training, they are often simplistic and may fail to capture the full complexity of the experimental conditions [51, 55, 14, 61].
@@ 3.1 Bragg disc detection and localization
Single-stage object detection models, such as YOLOv8, adopt a fully convolutional framework that directly predicts bounding box coordinates from feature maps [22, 54]. This anchor box-free design enables a well-generalized [14], object-center-aware feature augmentation, wherein grid cells serve as reference points for predicting offsets to object centers. Such a center-based approach is particularly well-suited for the high-throughput and high-precision detection of Bragg disc centers in electron diffraction patterns [42, 22, 8, 54].
@@
Where $c_{x}$ and $c_{y}$ are the column and row indices of the grid cell, $d_{n}$ is the bounding box parameter, and $s$ is the stride level.
To improve the center localization in our object detection model, we incorporated an efficient autocorrelation-based approach, leveraging the Wiener–Khinchin theorem [17, 57, 23]. Recently, the log-magnitude power spectrum of reciprocal-space intensities has been applied via cepstral analysis to decouple lattice periodicities in 4D-STEM diffraction patterns. Although the logarithmic operation compresses the dynamic range and suppresses dominant intensity variations, it can potentially obscure weaker signals in low SNR situations. In our implementation, spectral decomposition is performed without applying the logarithmic transform, similar to the Patterson function [40], utilizing an autocorrelation function to retain the full-intensity scaling. This strategy enhances the extraction of interatomic distances while minimizing potential ambiguities in Bragg disc intensities caused by sample tilt and thickness variations.
@@ 4.2 Strain mapping of pseudomorphic Si/SiGe epitaxial layers
To investigate the fidelity of our NDPD workflow in strain analysis, we acquired a 4D-STEM dataset from a Si/SiGe multilayer TEM sample at an operating voltage of 200 kV. The TEM foil thickness was approximately 70 nm, as determined by EELS. The acquisition conditions were optimized so that each diffraction disc spans at least 10 pixels in diameter and extends to at least the third-order Laue zone, ensuring both spatial precision and sufficient sampling of the diffraction space.
Fig. 6 illustrates a set of diffraction patterns from the Si substrate of the Si/SiGe TEM sample. The object detection model generates bounding box predictions, and from these predictions, we extract the center coordinates for each diffraction disc. The alternating Si and SiGe layers provide distinct light and dark contrast, as shown in the virtual annular dark-field STEM image in Fig. 7. Strain maps $\varepsilon_{xx}$ and $\varepsilon_{yy}$ represent the out-of-plane and in-plane strain fields populated from two orthogonal reciprocal lattice vectors, $\langle 100\rangle$ and $\langle 110\rangle$, respectively.
The strain calculations are conducted by tracking the center coordinates of diffraction discs by our object detection model and relative displacements measured with respect to a reference region selected in the strain-free Si substrate. The oscillations in the colored map of strain field $\varepsilon_{xx}$ indicate positive (tensile) strain along the growth direction. The profile is plotted along the $\langle 100\rangle$ direction, and the corresponding mean strain variation is approximately 0.8 %, which is in good agreement with the expected maximum theoretical value in SiGe layers.
@@
The strain variations due to the bending of the thin TEM foil are also evident in the strain field maps. These variations contribute to the strain modulations on the order of 0.1-0.2 % along the nominally strain-free $\langle 110\rangle$ direction in the $\varepsilon_{yy}$ strain map.
Computational benchmarking was conducted on a desktop system equipped with an Intel i7 8-core CPU and a 12GB NVIDIA RTX 3060 GPU. To evaluate the end-to-end latency of the strain analysis, we utilized a 4D-STEM dataset acquired from an ROI scanned over a 97$\times$40 pixels field of view. A total of 3880 diffraction patterns, each with an image resolution of 512$\times$512 pixels, were processed asynchronously in 35 seconds, yielding an effective throughput of 110 fps. The model inference was carried out at an input resolution of 512$\times$512 pixels per pattern.
Figure 7: Top left to right: the virtual DF image of Si/SiGe multilayers, the uniaxial strain map $\varepsilon_{xx}$ along the growth direction, and the orthogonal strain component $\varepsilon_{yy}$. Bottom left to right: the line profile across the thin-film growth $\langle 100\rangle$ and strain-free $\langle 110\rangle$ directions.
@@ 4.3 Phase and strain mapping of a multiphase metallic alloy
Figure 8: Top left to right: example electron diffraction patterns for matrix/coherent Phase I and incoherent Phases II and III overlapping with bcc matrix. bottom left to right: virtual DF image, and interatomic lattice displacement map.
Fig. 8 presents example displacement field maps derived from two orthogonal $\langle 110\rangle$ reciprocal lattice vectors, respectively. These maps reveal small, spherical bcc-precipitates (Phase I), approximately 50 nm in size, that are coherently embedded within the parent $\beta$-bcc matrix. In contrast, incoherent phases are identified as Phase II, corresponding to the ordered oxide $Ti_{3}O$, and Phase III, comprising hcp-precipitates that overlap with the parent bcc-matrix, as evidenced by the electron diffraction patterns.
The corresponding strain maps in Fig. 9 reveal a 1 % tensile (positive) strain $\varepsilon_{xx}$, accompanied by a 0.5 % compressive (negative) in $\varepsilon_{yy}$ and a 0.5 % shear (negative) strain in $\varepsilon_{xy}$, all originating from the embedded bcc-precipitates. In addition to strain components, the rotation map indicates an approximate $1^{\circ}$ lattice rotation in the coherency strain field. These findings support the implementation of strain engineering strategies that can be leveraged to improve the mechanical performance of the alloy.
Figure 9: Top left to right: strain maps of $\varepsilon_{xx}$ and $\varepsilon_{yy}$. Bottom left to right: strain map of $\varepsilon_{xy}$ and in-plane rotation map.
@@ 5 Conclusion
In this work, we introduce a parallelized post-processing framework for large-scale 4D-STEM datasets, leveraging an efficient neural network-based object detection model. Rapid and accurate object detection presents a promising avenue for the analysis of electron diffraction patterns and the localization of interatomic lattice distances. Our approach enables sub-pixel precision in the detection of diffraction features, enabling high-accuracy crystallographic analysis. The ability to rapidly and reliably extract crystallographic information from 4D-STEM datasets acquired under varying experimental conditions accelerates materials discovery, from semiconductor development to alloy design.
6 Data availability
Python code and a demonstration video for the neural diffraction pattern detection workflow are available at 4D-STEM-neural-diffraction-pattern-recognition-tempo4d.